Your company has an opportunity to invest in a project that is expected to result in after-tax cash flows of $13,000 the first year, $15,000 the second year, $18,000 the third year, -$8,000 the fourth year, $25,000 the fifth year, $31,000 the sixth year, $34,000 the seventh year, and -$6,000 the eighth year. The project would cost the firm $67,100. If the firm's cost of capital is 12%, what is the modified internal rate of return?
Answer #1

MIRR = ( FVc / PVfc )1/n -1

where, FVc is the  future value of positive cash flows

and PVfc is the present value of negative cash flows

CF0 = -67100
CF1 = 13000
CF2 = 15000
CF3 = 18000
CF4 = -8000
CF5 = 25000
CF6 = 31000
CF7 = 34000
CF8 = -6000

WACC = r = 12%

PVfc = ΣCFn/(1+r)n = 67100 + 8000/(1+0.12)4 + 6000/(1+0.12)8 = $74607.44

FVc = 13000(1+0.12)7 + 15000(1+0.12)6 + 18000(1+0.12)5 + 25000(1+0.12)3 + 31000(1+0.12)2 + 34000(1+0.12) = $202157.95

Hence, MIRR = (202157.95/74607.44)1/8 - 1 = 0.1327 = 13.27%